# PAIR-Former: Budgeted Relational Multi-Instance Learning for Functional miRNA Target Prediction

### v1

@@ 2 Related Work
A key practical challenge in miRNA targeting is that the CTS pool size is large and heavy-tailed across transcripts, making naive relational aggregation over all instances computationally infeasible. Unlike standard attention-based MIL or Set Transformer formulations that implicitly assume all instances can be encoded and attended to, BR-MIL enforces a hard per-bag budget $K$ on expensive encoding and relational computation. PAIR-Former instantiates this regime by combining (i) a cheap full-pool scan, (ii) budgeted and diversity-aware subset selection, and (iii) permutation-invariant Set Transformer aggregation on the selected tokens, while keeping candidate generation and alignment-aware CTS encoding TargetNet-compatible for protocol comparability (Min et al., 2022; Li and others, 2024). This design targets the gap between strong site scorers and realistic compute constraints in large-bag functional targeting.
@@ 3 Problem Setup
Permutation invariance requires $\hat{Y}(B)=\hat{Y}(\pi(B))$ for any permutation $\pi$ of the instances, which holds when selection depends only on the multiset of cheap signals (with fixed tie-breaking) and $f_{\phi}$ is permutation-invariant (e.g., Set Transformer / DeepSets). Per bag, the pool-dependent work is $\mathcal{O}(n)$ (cheap scan/selection) while the expensive relational component is governed by the budget, typically $\mathcal{O}(K)$ for expensive encoding and $\mathcal{O}(K^{2})$ for attention-based aggregation. Full notation, the formal BR-MIL definition/objective, and invariance conditions are deferred to Appendix A.
@@ 4.4 Budgeted subset selection with STSelector
Design principle.
Under heavy-tailed candidate pools, pure Top-$K$ selection tends to concentrate on near-duplicate sites. We therefore design STSelector to combine: (i) exploitation of high-confidence candidates under the cheap encoder, and (ii) coverage/diversity over transcript position and embedding space, while remaining CPU-only for low-latency inference.
This design is aligned with the approximation view in Sec. 5: by spreading the budget over high-scoring yet non-redundant candidates, the selector aims to cover a larger fraction of the total influence mass.
High-level behavior.
STSelector first selects a Top-$K_{1}$ subset according to cheap logits $\tilde{z}_{i}$, and then fills the remaining $K_{2}$ slots by encouraging coverage across transcript positions and reducing redundancy in the cheap embedding space. The output is a diverse subset $S\subseteq[n]$ with $|S|=K$ that preserves high-scoring candidates while avoiding over-concentration.
@@
Implementation details.
Full algorithmic steps (binning, SimHash-based deduplication, and balanced quota allocation) are provided in Appendix D.
@@ 4.7 Losses
The cheap encoder $\tilde{e}_{\tilde{\theta}}$ is trained to approximate the expensive teacher $e_{\theta}$ using a composite objective that combines: (i) supervised loss on student logits, (ii) temperature-scaled logit distillation, (iii) feature-level matching between teacher and student embeddings, and (iv) optional relational distillation via similarity matrices.
@@ 5.2 Generalization controlled by $K$
Full assumptions and proofs of Theorems 5.1 and 5.2 are provided in Appendix B.
@@ 6.3 Implementation details
All experiments are conducted on a single NVIDIA RTX 4090 GPU with 24 GB memory. We use PyTorch 2.4.1 (+cu121) with CUDA 12.1 and cuDNN 9.1.0.
@@ 6.4 Main results
Table 1: Overall performance on the released miRAWtest half-split (held-out test subsets). We report mean performance for reproduced baselines and our methods under the same split and the same evaluation code. Quoted baselines are taken from prior work under their original evaluation settings and are shown for context only (not directly comparable).
@@
Max pooling (ours, expensive encoder + max pooling) 0.8264 0.8125 0.8049 0.7822 0.8453 0.7646 0.8317
PAIR-Former (ours) 0.9961 0.9740 0.9744 0.9886 0.9598 0.9890 0.9612
@@
Under the released miRAWtest half-split and a fixed operating budget $K^{\star}{=}64$, PAIR-Former achieves strong performance (PR-AUC $0.9961$, F1@0.5 $0.9740$), outperforming TargetNet and a matched max-pooling baseline. We note that miRAWtest subsets are small and class-balanced, so absolute metrics may saturate; we focus on improvements under the identical split and evaluation code.
@@ 6.7 Selector ablation
The best selector variant depends on the budget. At K=32K{=}32, S2 improves over S0 by reducing redundancy via embedding deduplication. At K=64K{=}64, S1 is slightly better, suggesting that extra pruning can be overly aggressive when the budget is larger. We use S2 as the default due to its stronger small-budget performance and redundancy control in heavy-tailed pools.
@@ 7 Discussion
Our supervision is primarily pair-level; thus, learned instance importance/attention should not be interpreted mechanistically, and performance is bounded by label noise and curation artifacts. BR-MIL also inherits dependence on candidate generation: unretrieved functional sites are irrecoverable downstream. Hard selection under small budgets can miss weak-but-combinatorial evidence; more principled budgeted selection (e.g., stochastic/differentiable subset selection with exchangeable noise) is an important direction. Finally, Set Transformer attention scales as $\mathcal{O}(K^{2})$; sparse/linear attention or hierarchical pooling could extend BR-MIL to larger budgets and broader functional datasets beyond miRAW.
@@ 8 Conclusion
This paper presents work whose goal is to advance machine learning for budget-constrained relational multi-instance prediction, motivated by functional miRNA–mRNA target prediction. Potential positive impacts include improving the prioritization of candidate regulatory interactions for downstream experimental validation and enabling more compute-efficient analysis of large candidate pools. We do not anticipate major direct negative societal impacts of this work beyond standard risks of dataset bias and over-interpretation of model outputs.
@@ A.4 Computational Complexity
For each bag, cheap encoding costs $\mathcal{O}(n)$ and is cacheable; selection costs $\mathcal{O}(n)$ (up to additional diversity heuristics). Expensive encoding costs $\mathcal{O}(K)$ and relational aggregation typically costs $\mathcal{O}(K^{2})$. Thus the deployed cost per bag is
$\mathcal{O}(n)+\mathcal{O}(K^{2}),$
under a strict access budget $K$, versus $\mathcal{O}(n^{2})$ if all instances were processed relationally.
@@ B.1 Preliminaries: masking view of budgeted execution
B.3 Generalization controlled by the budget $K$ (Theorem 5.2)

### v2

@@ 2 Related Work
Budgeted MIL in computational pathology.
Gigapixel whole-slide image (WSI) classification faces a similar computational challenge: exhaustive patch-level encoding is prohibitive. Methods such as Campanella et al. (Campanella et al., 2019), CLAM (Lu et al., 2021), ZoomMIL (Thandiackal et al., 2022), DTFD-MIL (Zhang et al., 2022), MagNet (Dimitriou and Arandjelović, 2024), and TransMIL (Shao and others, 2021) address this via iterative selection, clustering-constrained attention, differentiable zooming, or pseudo-bag decomposition. While these share the high-level scan-select-aggregate structure with BR-MIL, they typically rely on heuristic or scale-driven selection without explicit diversity constraints or formal approximation guarantees, and are tailored to spatial patch structures rather than heavy-tailed sequence-derived candidate pools.
@@
Beyond computational pathology, a key practical challenge in miRNA targeting is that the CTS pool size is large and heavy-tailed across transcripts, making naive relational aggregation over all instances computationally infeasible. Unlike standard attention-based MIL or Set Transformer formulations that implicitly assume all instances can be encoded and attended to, BR-MIL enforces a hard per-bag budget $K$ on expensive encoding and relational computation. PAIR-Former instantiates this regime by combining (i) a cheap full-pool scan, (ii) budgeted and diversity-aware subset selection, and (iii) permutation-invariant Set Transformer aggregation on the selected tokens, while keeping candidate generation and alignment-aware CTS encoding TargetNet-compatible for protocol comparability (Min et al., 2022; Li and others, 2024). This design targets the gap between strong site scorers and realistic compute constraints in large-bag functional targeting.
@@ 3 Problem Setup
Permutation invariance requires $\hat{Y}(B)=\hat{Y}(\pi(B))$ for any permutation $\pi$ of the instances, which holds when selection depends only on the multiset of cheap signals (with fixed tie-breaking) and $f_{\phi}$ is permutation-invariant (e.g., Set Transformer / DeepSets). Per bag, the computational challenge decomposes into two distinct bottlenecks: (a) per-instance encoding cost—each CTS requires an expensive forward pass through $e_{\theta}$, so naively encoding all $n$ candidates costs $\mathcal{O}(n)$ expensive forward passes, which is prohibitive under heavy-tailed $n$ (see Section 6.2 for empirical distribution); and (b) relational aggregation cost—attention-based set functions scale as $\mathcal{O}(n^{2})$ in the number of encoded instances, which becomes intractable for large bags even if encoding were free. BR-MIL addresses both by restricting expensive encoding to $K$ selected instances ($\mathcal{O}(K)$ encoding cost) and performing relational aggregation only on the $K$-element subset ($\mathcal{O}(K^{2})$ aggregation cost), while a cheap $\mathcal{O}(n)$ scan covers the full pool. Full notation, the formal BR-MIL definition/objective, and invariance conditions are deferred to Appendix A.
@@ 4.4 Budgeted subset selection with STSelector
Motivation and design principle.
In miRNA targeting, functionally relevant CTSs are often spread across different regions of the 3'UTR, and sites with similar sequences or nearby positions tend to yield similar expensive representations. Pure Top-$K$ selection by cheap score may allocate budget to near-redundant sites in the same region, particularly in heavy-tailed pools where $n$ can exceed 10,000. STSelector addresses this via position binning and embedding deduplication, providing a compute-efficient, CPU-only selection mechanism with interpretable diversity constraints. Formally, this corresponds to improving coverage of the influence mass $\alpha_{i}$ in Theorem 5.1. The design combines (i) exploitation of high-confidence candidates and (ii) coverage across transcript position bins and embedding-space neighborhoods (via SimHash dedup). We note that accuracy gains from diversity-aware selection are more pronounced at smaller budgets ($K{\leq}32$) where redundancy is more costly; at K=64K{=}64 on saturated benchmarks, differences from TopK are modest (Table 3).
Selection procedure.
STSelector operates in three stages: (1) Top-$K_{1}$ exploitation: select the $K_{1}$ candidates with highest cheap logits $\tilde{z}_{i}$. (2) Position-binned diversity pool: partition the remaining candidates into $B$ bins by normalized transcript position, maintain a per-bin heap of the top-$m$ candidates by $\tilde{z}_{i}$, and remove near-duplicates via SimHash on $\tilde{h}_{i}$. (3) Balanced quota allocation: assign the $K_{2}$ slots across bins proportionally to each bin’s aggregate cheap confidence (with a minimum of one per bin), yielding a diverse subset $S_{2}$. The final output is $S=\mathrm{dedup}(S_{1}\cup S_{2})$, padded to $|S|=K$ by descending $\tilde{z}_{i}$ if needed.
@@
Full algorithm.
Pseudocode and hyperparameter defaults ($B,m,c,\tau_{w}$) are provided in Appendix D.
@@ 4.7 Losses
The cheap encoder $\tilde{e}_{\tilde{\theta}}$ is trained to approximate the expensive teacher $e_{\theta}$ using a composite objective that combines: (i) supervised loss on student logits, (ii) temperature-scaled logit distillation, (iii) feature-level matching between teacher and student embeddings, and (iv) optional relational distillation via similarity matrices (shown to provide marginal gains at the current scale; see Appendix E).
@@ 5.2 Generalization controlled by $K$
Full assumptions and proofs of Theorems 5.1 and 5.2 are provided in Appendix B. Notably, the Rademacher complexity bound (formerly stated as an assumption) is now derived from the Set Transformer architecture in Appendix B.3. Appendix B.5 provides practical guidance for budget selection based on the bias-variance tradeoff, and Appendix B.6 discusses the tightness of the $\sqrt{K}$ bound via a matching lower bound.
@@ 6.1 Datasets and evaluation protocol
Ground-truth source.
The miRAWtest positive pairs are derived from experimentally validated miRNA–mRNA interactions curated from multiple sources, including CLASH and PAR-CLIP assays that directly capture Argonaute-mediated miRNA–target duplexes, as well as validated functional targets from Diana TarBase and MirTarBase (Pla and others, 2018; Helwak et al., 2013; Hafner et al., 2010). Negative pairs are constructed from experimentally confirmed non-functional interactions rather than synthetic sequences, ensuring that the benchmark reflects genuine biological targeting mechanisms (Pla and others, 2018). This makes miRAWtest one of the few miRNA target benchmarks with experimentally grounded functional labels, albeit limited in scale.
@@
Data partitioning and supervision granularity.
Stage 1–2 use the miRAW Train-Validation split for CTS-level supervision/distillation on local 40nt windows, while Stage 3 uses a half-split of miRAWtest for pair-level training and evaluation. Specifically, Stage-3 training uses subsets $\{1,\dots,5\}$ (with a 0.9/0.1 train-validation split), and held-out testing uses subsets $\{0,6,\dots,9\}$. Although some transcript identities recur across Stage 1–2 and Stage-3, the supervision differs fundamentally: Stage 1–2 learn local CTS functionality, whereas Stage 3 predicts transcript-level functional repression for full miRNA–mRNA pairs.
@@ 6.4 Implementation details
All experiments are conducted on a single NVIDIA RTX 4090 GPU with 24 GB memory for reproducibility and accessibility; the pipeline contains no architectural barriers to multi-GPU data-parallel training.
@@ 6.5 Main results
[…]
@@
[…]
@@
Table 2: External validation on deepTargetPro benchmark (10-fold cross-validation). We use splits 1–5 for training and splits 0,6–9 for testing (no overlap). “PAIR-Former (transfer)” uses miRAW-pretrained CTS encoders and trains only the Stage 3 aggregator on deepTargetPro. “PAIR-Former (full)” trains all three stages on deepTargetPro from scratch. Baseline results are quoted from TEC-miTarget (Yang et al., 2024) (Table 7–8) on all 10 test splits. All metrics are reported as percentages.
Method Acc (%) Sens (%) Spec (%) PPV (%) NPV (%) F1 (%)
Seed-match-based methods (quoted from TEC-miTarget)
PITA (Kertesz et al., 2007) 50.53 13.65 87.41 51.96 50.31 21.62
mirSVR (Betel et al., 2010) 50.01 27.76 72.26 49.97 50.01 35.68
miRDB (Wong and Wang, 2015) 53.73 12.39 95.07 71.35 52.05 21.10
microT (Paraskevopoulou et al., 2013) 61.13 58.94 63.32 61.62 60.70 60.24
TargetScan (Lewis et al., 2005) 55.77 39.45 72.08 58.52 54.36 47.12
Deep learning methods (quoted from TEC-miTarget)
deepTarget (Wen and others, 2016) 65.21 34.77 93.54 83.32 60.64 49.04
deepTargetPro (Lee et al., 2020) 78.04 75.51 80.38 78.17 77.92 76.81
TargetNet (Min et al., 2022) 72.61 95.08 51.67 64.69 91.90 76.99
TEC-miTarget (Yang et al., 2024) 79.97 78.56 81.29 79.67 80.25 79.11
Ours (5 test splits: 0,6–9)
[…]
@@
Under the released miRAWtest half-split and a fixed operating budget $K^{\star}{=}64$, PAIR-Former achieves strong performance (PR-AUC $0.989{\pm}0.001$, F1@0.5 $0.975{\pm}0.002$), outperforming TargetNet and a matched max-pooling baseline. We note that miRAWtest subsets are small and class-balanced, so absolute metrics may saturate; we focus on improvements under the identical split and evaluation code. We additionally verified robustness to the choice of train/test partition by evaluating on two alternative splits (consecutive and interleaved); performance remains consistent across all three partitioning strategies (PR-AUC $0.993{\pm}0.003$ over 3 splits $\times$ 3 seeds; see Appendix I).
External validation on deepTargetPro.
[…]
@@ 6.8 Selector ablation
At K=64K{=}64, selector variants show small differences (${\sim}0.0003$ PR-AUC), reflecting benchmark saturation at PR-AUC ${\sim}0.994$. At K=32K{=}32, S2 shows slightly stronger performance, suggesting diversity-aware selection may provide more benefit when the budget is tighter and redundancy is more costly. We use S2 as the default due to its principled diversity constraints (position binning, embedding deduplication) and compute-efficiency (CPU-only, low overhead), which may be valuable for heavy-tailed pools even when accuracy gains are modest at large $K$ on saturated benchmarks.
@@ 7 Discussion
[…]
@@ 8 Conclusion
This paper presents work whose goal is to advance machine learning for budget-constrained relational multi-instance prediction, motivated by functional miRNA–mRNA target prediction. Potential positive impacts include improving the prioritization of candidate regulatory interactions for downstream experimental validation, enabling more compute-efficient analysis of large candidate pools, and accelerating the development of RNA-based therapeutics (e.g., antagomirs and miRNA mimics) for cancer and genetic diseases. However, the same prediction capabilities could, in principle, be leveraged to design interventions that deliberately disrupt normal gene regulatory networks or to identify and exploit tissue-specific regulatory vulnerabilities—concerns that are relevant to the broader field of precision biology. We note that computational predictions of miRNA targeting require experimental validation before any clinical or therapeutic application, and we encourage responsible use. Beyond these domain-specific considerations, standard risks apply, including dataset bias, population underrepresentation in training data, and over-interpretation of model outputs.
@@ A.4 Computational Complexity
For each bag, the computational cost decomposes into two independent bottlenecks:
(a) Per-instance encoding cost. Running the expensive encoder $e_{\theta}$ on all $n$ candidates would require $\mathcal{O}(n)$ expensive forward passes. For heavy-tailed $n$ (e.g., $n>500$; see Figure 2 for empirical distribution), this becomes prohibitive even if aggregation were free. BR-MIL avoids this by using a cheap encoder $\tilde{e}_{\tilde{\theta}}$ to scan all $n$ candidates in $\mathcal{O}(n)$ cheap operations, then applying $e_{\theta}$ only to the selected $K$ instances, reducing expensive encoding to $\mathcal{O}(K)$.
(b) Relational aggregation cost. Attention-based aggregators (e.g., Set Transformer) scale as $\mathcal{O}(n^{2})$ in the number of encoded instances due to pairwise interactions. Even if encoding were free, $\mathcal{O}(n^{2})$ aggregation is intractable for large $n$. BR-MIL restricts aggregation to the $K$-element selected subset, reducing this to $\mathcal{O}(K^{2})$.
Thus the deployed cost per bag is
$\mathcal{O}(n)_{\text{cheap scan}}+\mathcal{O}(K)_{\text{expensive encoding}}+\mathcal{O}(K^{2})_{\text{aggregation}},$
under a strict access budget $K$, versus $\mathcal{O}(n)_{\text{expensive encoding}}+\mathcal{O}(n^{2})_{\text{aggregation}}$ if all instances were processed relationally.
@@ B.1 Preliminaries: masking view of budgeted execution
B.3 Derivation of Rademacher Complexity Bound (Former Assumption B.24)
In this section, we derive the Rademacher complexity bound for the Set Transformer aggregator, which was previously stated as Assumption B.24. We show that this bound follows from the specific architecture of the Set Transformer and standard generalization theory for deep neural networks (Bartlett et al., 2017).
Roadmap.
We proceed in two steps: (i) bound the spectral norms of the Set Transformer’s parameter matrices (Sec. B.3.1), (ii) apply the spectral norm-based Rademacher complexity bound from Bartlett et al. (2017) (Sec. B.3.2).
B.3.1 Step 1: Spectral Norm Bounds for Set Transformer
We first recall the Set Transformer architecture and establish spectral norm bounds for its components.
Architecture recap.
The Set Transformer $f_{\phi}$ operates on $K$ tokens $X\in\mathbb{R}^{K\times d_{z}}$ and consists of: $L$ Self-Attention Blocks (SAB), each containing multi-head attention (MHA) with projection matrices $W_{Q},W_{K},W_{V},W_{O}$, a feed-forward network (FFN) with weight matrices $W_{1},W_{2}$, layer normalization (LN) with learnable parameters $\gamma,\beta$, and residual connections; followed by Pooling by Multi-head Attention (PMA) with learnable seed vectors, and a final linear projection $W_{\text{out}}$ to scalar output.
Assumption B.9 (Bounded spectral norms).
All parameter matrices in the Set Transformer have spectral norms bounded by a constant $\sigma_{\max}$: $\|W_{Q}\|_{2},\|W_{K}\|_{2},\|W_{V}\|_{2},\|W_{O}\|_{2},\|W_{1}\|_{2},\|W_{2}\|_{2},\|W_{\text{out}}\|_{2}\leq\sigma_{\max}$. Layer normalization parameters satisfy $\|\gamma\|_{\infty},\|\beta\|_{\infty}\leq 1$. This is mild and holds in practice under standard initialization (e.g., Xavier/He), spectral normalization (Miyato et al., 2018), or weight decay regularization.
Lemma B.10 (Product of spectral norms).
Under Assumption B.9, the product of spectral norms across all layers of the Set Transformer is bounded by $\prod_{i=1}^{N_{\text{layers}}}\|W_{i}\|_{2}\leq\sigma_{\max}^{N_{\text{layers}}}$, where $N_{\text{layers}}=6L+O(1)$ is the total number of weight matrices (4 for MHA and 2 for FFN per SAB layer, plus PMA and final projection).
Proof.
Each SAB layer contains 6 weight matrices. The product of spectral norms is at most $\sigma_{\max}^{N_{\text{layers}}}$. ∎
Lemma B.11 (Sum of squared Frobenius-to-spectral ratios).
Under Assumption B.9, the sum of squared Frobenius-to-spectral ratios satisfies $\sum_{i=1}^{N_{\text{layers}}}\|W_{i}\|_{F}^{2}/\|W_{i}\|_{2}^{2}\leq N_{\text{layers}}\cdot d_{\max}$, where $d_{\max}=\max(d_{\text{model}},d_{\text{ff}})$.
Proof.
For any $W\in\mathbb{R}^{m\times n}$, $\|W\|_{F}^{2}\leq\min(m,n)\cdot\|W\|_{2}^{2}$. Summing over all layers yields the bound. ∎
B.3.2 Step 2: Rademacher Complexity via Spectral Norm Bounds
We now apply the spectral norm-based Rademacher complexity bound from Bartlett et al. (2017).
Theorem B.12 (Rademacher complexity for deep networks – Bartlett et al. (2017)).
Let $\mathcal{F}$ be a class of $L$-layer neural networks with weight matrices $W_{1},\ldots,W_{L}$, 1-Lipschitz activation functions (e.g., ReLU, softmax), and inputs bounded by $\|x\|_{2}\leq R_{\text{in}}$. Then the Rademacher complexity on $M$ samples satisfies
$\mathfrak{R}_{M}(\mathcal{F})\leq\frac{C}{\sqrt{M}}\cdot\Bigg(\prod_{i=1}^{L}\|W_{i}\|_{2}\Bigg)\cdot\sqrt{\sum_{i=1}^{L}\frac{\|W_{i}\|_{F}^{2}}{\|W_{i}\|_{2}^{2}}}\cdot R_{\text{in}},$ (18)
where $C=O(\sqrt{L})$ is a universal constant. Crucially, this bound depends on the input norm $R_{\text{in}}$ rather than the input dimension, which is the key to obtaining $\sqrt{K}$ instead of $K$.
Theorem B.13 (Rademacher complexity of Set Transformer – main result).
Let $\mathcal{F}_{K}$ be the class of Set Transformer aggregators satisfying Assumption B.9, operating on at most $K$ tokens of dimension $d_{z}$, with bounded token norms $\|z_{i}\|_{2}\leq R$ for all $i$. Then the Rademacher complexity satisfies
$\mathfrak{R}_{M}(\mathcal{F}_{K})\leq C_{\text{arch}}\cdot\frac{R\sqrt{K}}{\sqrt{M}},$ (19)
where $C_{\text{arch}}$ depends on $(L,d_{\text{model}},d_{\text{ff}},\sigma_{\max})$ but not on $K$ or $n$.
Proof.
We apply Theorem B.12 to the Set Transformer. The input matrix $X\in\mathbb{R}^{K\times d_{z}}$ with $\|z_{i}\|_{2}\leq R$ satisfies $\|X\|_{F}=\sqrt{\sum_{i=1}^{K}\|z_{i}\|_{2}^{2}}\leq R\sqrt{K}$, so $R_{\text{in}}=R\sqrt{K}$. By Lemmas B.10B.11, the spectral norm product is $\sigma_{\max}^{O(L)}$ and the Frobenius sum is $\sqrt{O(L\cdot d_{\max})}$. Substituting into (18) yields $\mathfrak{R}_{M}(\mathcal{F}_{K})\leq C\cdot\sigma_{\max}^{O(L)}\cdot\sqrt{O(L\cdot d_{\max})}\cdot R\sqrt{K}/\sqrt{M}$, which simplifies to (19) with $C_{\text{arch}}$ absorbing all architecture-dependent constants. ∎
Remark B.14 (Why $\sqrt{K}$ and not $K$?).
The $\sqrt{K}$ dependence arises solely from the input norm $\|X\|_{F}\leq R\sqrt{K}$. The key insight is that Theorem B.12 depends on the input norm $R_{\text{in}}$, not the input dimension $K\cdot d_{z}$. The Set Transformer has a fixed number of parameters (independent of $K$), so its capacity is controlled by parameter norms, not by $K$. The only way $K$ enters is through $\|X\|_{F}\leq R\sqrt{K}$, which scales as $\sqrt{K}$ due to the $\ell_{2}$ geometry of the input space. A naive Lipschitz-based bound would incorrectly give $O(K/\sqrt{M})$ by treating the input dimension as dominant.
Remark B.15 (Independence from bag size $n$).
The bound (19) depends only on $K$ (the budget) and not on $n$ because: (i) the selector $S$ is fixed (or conditioned upon); (ii) the Set Transformer only sees $K$ tokens, regardless of $n$; (iii) the number of parameters is independent of both $K$ and $n$. This validates the BR-MIL design: by budgeting expensive computation to $K$ tokens, we control generalization via $K$ rather than $n$.
Conclusion.
We have derived the Rademacher complexity bound (formerly Assumption B.24) from the Set Transformer architecture. The key steps were: (i) establishing spectral norm bounds (Assumption B.9, Lemmas B.10B.11), (ii) applying Bartlett et al.’s bound (Theorem B.12), (iii) observing that $\sqrt{K}$ arises from the input norm $R\sqrt{K}$ (Theorem B.13, Remark B.14). This bound is now a theorem rather than an assumption, strengthening the theoretical foundation of Theorem 5.2.
B.4 Generalization controlled by the budget $K$ (Theorem 5.2)